# A Differentiable Alignment Framework for Sequence-to-Sequence Modeling via Optimal Transport

### v1

@@ Abstract
Accurate sequence-to-sequence (seq2seq) alignment is critical for applications like medical speech analysis and language learning tools relying on automatic speech recognition (ASR). State-of-the-art end-to-end (E2E) ASR systems, such as the Connectionist Temporal Classification (CTC) and transducer-based models, suffer from peaky behavior and alignment inaccuracies. In this paper, we propose a novel differentiable alignment framework based on one-dimensional optimal transport, enabling the model to learn a single alignment and perform ASR in an E2E manner. We introduce a pseudo-metric, called Sequence Optimal Transport Distance (SOTD), over the sequence space and discuss its theoretical properties. Based on the SOTD, we propose Optimal Temporal Transport Classification (OTTC) loss for ASR and contrast its behavior with CTC. Experimental results on the TIMIT, AMI, and LibriSpeech datasets show that our method considerably improves alignment performance, though with a trade-off in ASR performance when compared to CTC. We believe this work opens new avenues for seq2seq alignment research, providing a solid foundation for further exploration and development within the community.
Optimal Transport, Sequence to Sequence, Alignment, Speech, ASR
@@ 1 Introduction
Sequence-to-sequence (seq2seq) alignment is a fundamental challenge in automatic speech recognition (ASR) systems, where, beyond text prediction, precise alignment of text to the corresponding speech is crucial for many applications. For example, in the medical domain, accurate alignment helps speech and language pathologists pinpoint specific speech segments for analyzing pathological cues, such as stuttering or voice disorders. In real-time subtitling, precise alignment ensures that subtitles are synchronized with spoken words, which is crucial for live broadcasts and streaming content. In language learning tools, ASR systems use alignment to provide feedback on pronunciation and fluency, allowing learners to compare their speech to target pronunciations. In these ASR-driven applications, while word error rate (WER) is an important performance metric, frame-level and word-level alignment accuracy are equally important for improving the system’s applicability and responsiveness.
@@
In the literature, two primary approaches to ASR have emerged, i.e., hybrid systems and end-to-end (E2E) models. In hybrid approaches, a deep neural network-hidden Markov model (DNN-HMM) (Morgan & Bourlard, 1990; Bourlard & Morgan, 2012; Young, 1996; Povey, 2005; Abdel-Hamid et al., 2012; Graves et al., 2013a; Dahl et al., 2012) system is typically trained, where the DNN is optimized by minimizing cross-entropy loss on the forced alignments generated for each frame of audio embeddings from a hidden Markov model-Gaussian mixture model (HMM-GMM). One notable disadvantage of the hybrid approach is that the model cannot be optimized in an E2E manner, which may result in suboptimal performance (Hannun, 2014). More recently, E2E models for ASR have become very popular due to their superior performance. There are three popular approaches for training an E2E model: (i) attention-based encoder-decoder (AED) models (Chan et al., 2015; Radford et al., 2023; Watanabe et al., 2017; Prabhavalkar et al., 2023), (ii) using Connectionist Temporal Classification (CTC) loss (Graves et al., 2006; Graves & Jaitly, 2014), and (iii) neural Transducer-based models (Graves, 2012; Kuang et al., 2022; Graves et al., 2013b). AED models use an encoder to convert the input audio sequence into a hidden representation. The decoder, typically auto-regressive, generates the output text sequence by attending to specific parts of the input through an attention mechanism, often referred to as soft alignment (Yan et al., 2022) between the audio and text sequences. This design, however, can make it challenging to obtain word-level timestamps and to do teacher-student training with soft labels. Training AED models also requires a comparatively large amount of data, which can be prohibitive in low-resource setups. In contrast to AED models, CTC and transducer-based models maximize the marginal probability of the correct sequence of tokens (transcript) over all possible valid alignments (paths), often referred to as hard alignment (Yan et al., 2022). However, recent research has shown that only a few paths, which are dominated by blank labels, contribute meaningfully to the marginalization, leading to the well-known peaky behavior that can result in suboptimal ASR performance (Zeyer et al., 2021). Unfortunately, it is not possible to directly identify these prominent paths, or those that do not disproportionately favor blank labels, in advance within E2E models. This observation serves as the main motivation of our work.
@@
  • We propose a novel, parameterized, and differentiable sequence-to-sequence alignment model with linear complexity both in time and space.
@@
  • Finally, we conduct proof-of-concept experiments on the TIMIT (Garofolo et al., 1993), AMI (Carletta et al., 2005), and Librispeech (Panayotov et al., 2015) datasets, demonstrating that our method mitigates the peaky beahavior, improves alignment performance, and achieves promising results in E2E ASR.
@@ 2 Related Work
CTC loss. The CTC criterion (Graves et al., 2006) is a versatile method for learning alignments between sequences. This versatility has led to its application across various sequence-to-sequence tasks (Liu et al., 2020; Chuang et al., 2021; Yan et al., 2022; Gu & Kong, 2021; Graves & Schmidhuber, 2008; Molchanov et al., 2016). However, despite its widespread use, CTC has numerous limitations that impact its effectiveness in real-world applications. To address issues such as peaky behavior (Zeyer et al., 2021), label delay (Tian et al., 2023), and alignment drift (Sak et al., 2015), researchers have proposed various extensions. These extensions aim to refine the alignment process, ensuring better performance across diverse tasks. Delay-penalized CTC (Yao et al., 2023) and blank symbol regularization (Yang et al., 2023; Zhao & Bell, 2022; Bluche et al., 2015) attempt to mitigate label delay issues. Other works have tried to control alignment through teacher model spikes (Ghorbani et al., 2018; Kurata & Audhkhasi, 2019) or external supervision (Zeyer et al., 2020; Senior et al., 2015; Plantinga & Fosler-Lussier, 2019), though this increases complexity. Recent advancements like Bayes Risk CTC offer customizable, end-to-end approaches to improve alignment without relying on external supervision (Tian et al., 2023).
Transducer loss. The transducer loss was introduced to address the conditional independence assumption of CTC by incorporating a predictor network (Graves, 2012). However, similarly to CTC, transducer models suffer from label delay and peaky behavior (Yu et al., 2021). To mitigate these issues, several methods have been proposed, such as e.g., Pruned RNN-T (Kuang et al., 2022), which prunes alignment paths before loss computation, FastEmit (Yu et al., 2021), which encourages faster symbol emission, delay-penalized transducers (Kang et al., 2023), which add a constant delay to all non-blank log-probabilities, and minimum latency training (Shinohara & Watanabe, 2022), which augments the transducer loss with the expected latency. Further extensions include CIFTransducer (CIF-T) for efficient alignment (Zhang et al., 2024), self-alignment techniques (Kim et al., 2021), and lightweight transducer models using CTC forced alignments (Wan et al., 2024).
@@ 3 Problem Formulation
We define $\mathcal{U}_{\leq N}^{d}=\bigcup_{n\leq N}\mathcal{U}_{n}^{d}$ to be the set of all $d$-dimensional vector sequences of length at most $N$. Let us consider a distribution $\mathcal{D}_{\mathcal{U}_{\leq N}^{d}\times\mathcal{U}_{\leq N}^{d}}$ and pairs of sequences ($(\{{\bm{x}}_{i}\}_{i=1}^{n},\{{\bm{y}}_{i}\}_{i=1}^{m})$ of length $n$ and $m$ drawn from $\mathcal{D}_{\mathcal{U}_{\leq N}^{d}\times\mathcal{U}_{\leq N}^{d}}$. For notational simplicity, the sequences of the pairs ($(\{{\bm{x}}_{i}\}_{i=1}^{n},\{{\bm{y}}_{i}\}_{i=1}^{m})$ will be respectively denoted by $\{{\bm{x}}\}_{n}$ and $\{{\bm{y}}\}_{m}$ in the following. The goal in sequence-to-sequence tasks is to train a classifier that can accurately predict the target sequence $\{{\bm{y}}\}_{m}$ from the input sequence $\{{\bm{x}}\}_{n}$, enabling it to generalize to unseen examples. Typically, $n\neq m$, creating challenges for accurate prediction as there is no natural alignment between the two sequences. In this paper, we introduce a framework to address this class of problems, applying it specifically to the ASR domain. In this context, the first sequence $\{{\bm{x}}\}_{n}$ represents an audio signal, where each vector ${\bm{x}}_{i}\in\mathbb{R}^{d}$ corresponds to a time frame in the acoustic embedding space. The second sequence $\{{\bm{y}}\}_{m}$ is the textual transcription of the audio, where each element ${\bm{y}}_{i}$ belongs to a predefined vocabulary $L=\{l_{1},\dots,l_{|L|}\}$, such that $\{{\bm{y}}\}_{m}\in L^{m}$, where $L^{m}$ denotes the set of all $m$-length sequences formed from the vocabulary $L$.
Figure 1: Example of an alignment between embeddings of frames and target sequence. Red bullets represent the elements of the target sequence $\{{\bm{y}}\}_{m}$, while the blue bullets indicate the frame embeddings $\{{\bm{x}}\}_{n}$. In OTTC, the alignment guides the prediction model $F$ in determining which frames should map to which labels. Additionally, the alignment model has the flexibility to leave some frames unaligned, as represented by the blue-and-white bullets, allowing those frames to be dropped during inference.
@@ 4 Optimal Temporal Transport Classification
The core idea is to model the alignment between two sequences as a mapping to be learned along with the frame labels (see Figure 1). As the classification of audio frames improves, inferring the correct alignment becomes easier. Conversely, accurate alignments also improve frame classification. This mutual reinforcement between alignment and classification highlights the benefit of addressing both tasks simultaneously, contrasting with traditional hybrid models that treat them as separate tasks (Morgan & Bourlard, 1990). To achieve this, we propose the Sequence Optimal Transport Distance (SOTD), a framework for constructing pseudo-metrics over the sequence space $\mathcal{U}_{\leq N}^{d}$, based on a differentiable, parameterized model that learns to align sequences. Using this framework, we derive the Optimal Temporal Transport Classification (OTTC) loss, which allows the model to learn both the alignment and the classification in a unified manner.
@@ 4.2 Differentiable Temporal Alignment with Optimal Transport
In the following, we introduce 1D OT and define our alignment model. Consider the 1D discrete distributions $\mu[\bm{\alpha},n]$ and $\nu[\bm{\beta},m]$ expressed as superpositions of $\delta$ measures, i.e., a distribution that is zero everywhere except at a single point, where it integrates to 1:
@@
The bins of $\mu[\bm{\alpha},n]$ and $\nu[\bm{\beta},m]$ are $\llbracket 1,n\rrbracket$ and $\llbracket 1,m\rrbracket$, respectively, whereas the weights $\alpha_{i}$ and $\beta_{i}$ are components of the vectors $\bm{\alpha}\in\Delta^{n}$ and $\bm{\beta}\in\Delta^{m}$, with $\Delta^{n}$ the simplex set defined as $\Delta^{n}=\{\mathbf{v}\in\mathbb{R}^{n}|{0\leq v_{i}\leq 1,\sum_{i=1}^{n}v_{i}=1}\}\subset\mathbb{R}^{n}$. OT theory provides an elegant and versatile framework for computing distances between distributions such as $\mu[\bm{\alpha},n]$ and $\nu[\bm{\beta},m]$, depending on the choice of the cost function (Peyré & Cuturi, 2019) (chapter 2.4). One such distance is the 2-Wasserstein distance $\mathcal{W}_{2}$, which measures the minimal cost of transporting the weight of one distribution to match the other. This distance is defined as
$\mathcal{W}_{2}(\mu[\bm{\alpha},n],\nu[\bm{\beta},m])=\min_{\bm{\gamma}\in\Gamma^{\bm{\alpha},\bm{\beta}}}\sum_{i,j=1}^{n,m}\gamma_{i,j}|\!|i-j|\!|_{2}^{2},$ (2)
were $|\!|i-j|\!|_{2}^{2}$ is the cost of moving weight from bin $i$ to bin $j$ and $\gamma_{i,j}$ is the amount of mass moved from $i$ to $j$. The optimal coupling matrix $\bm{\gamma}^{*}$ is searched within the set of valid couplings $\Gamma^{\bm{\alpha},\bm{\beta}}$ defined as
@@
This constraint ensures that the coupling conserves mass, accurately redistributing all weights between the bins. A key property of optimal transport in 1D is its monotonicity (Peyré, 2019). Specifically, if there is mass transfer between bins $i$ and $j$ (i.e., $\gamma_{i,j}^{*}>0$) and similarly between bins $k$ and $l$ (i.e., $\gamma_{k,l}^{*}>0$), then it must hold that $i\leq k\Rightarrow j\leq l$. Consequently, when $\bm{\beta}$ has no zero components – meaning that every bin from $\nu$ is reached by the transport – the set $\{(i,j)\in[\!|1,n|\!]\times[\!|1,m|\!]\ |\ \bm{\gamma}_{i,j}^{*}>0\}$ satisfies the conditions of Definition 1, thereby forming a discrete monotonic alignment. This demonstrates that the optimal coupling can effectively model such alignments (see Figure 2).
Note: In the 1D case, the solution $\bm{\gamma}^{*}$ is unique and depends only on the number of distinct bins and their weights, not their specific values. Thus, the choice of $\llbracket 1,n\rrbracket$ and $\llbracket 1,m\rrbracket$ as bins is arbitrary (Peyré, 2019).
@@
Lemma 1: The function $\bm{\alpha}\mapsto\bm{\gamma}_{n}^{m,\bm{\beta}}(\bm{\alpha})$ is bijective from $\Delta^{n}$ to $\Gamma^{*,\bm{\beta}}[n]$
@@
4.2.1 Sequences-to-Sequences Distance

### v2

@@ Abstract
Accurate sequence-to-sequence (seq2seq) alignment is critical for applications like medical speech analysis and language learning tools relying on automatic speech recognition (ASR). State-of-the-art end-to-end (E2E) ASR systems, such as the Connectionist Temporal Classification (CTC) and transducer-based models, suffer from peaky behavior and alignment inaccuracies. In this paper, we propose a novel differentiable alignment framework based on one-dimensional optimal transport, enabling the model to learn a single alignment and perform ASR in an E2E manner. We introduce a pseudo-metric, called Sequence Optimal Transport Distance (SOTD), over the sequence space and discuss its theoretical properties. Based on the SOTD, we propose Optimal Temporal Transport Classification (OTTC) loss for ASR and contrast its behavior with CTC. Experimental results on the TIMIT, AMI, and LibriSpeech datasets show that our method considerably improves alignment performance compared to CTC and the more recently proposed Consistency-Regularized CTC, though with a trade-off in ASR performance. We believe this work opens new avenues for seq2seq alignment research, providing a solid foundation for further exploration and development within the community.
@@ 1 Introduction
Sequence-to-sequence (seq2seq) alignment is a fundamental challenge in automatic speech recognition (ASR), where, beyond text prediction, precise alignment of text to the corresponding speech is crucial for many applications. For example, in medical domain, accurate alignment helps speech and language pathologists pinpoint speech segments for analyzing pathological cues, such as stuttering or voice disorders. In real-time subtitling, precise alignment ensures that subtitles are synchronized with spoken words, which is crucial for live broadcasts and streaming content. In language learning tools, ASR systems use alignment to provide feedback on pronunciation and fluency, allowing learners to compare their speech to target pronunciations. In these ASR-driven applications, while word error rate (WER) is an important performance metric, frame-level and word-level alignment accuracy are equally important for improving the system’s applicability and responsiveness.
@@
In the literature, two primary approaches to ASR have emerged, i.e., hybrid systems and end-to-end (E2E) models. In hybrid approaches, a deep neural network-hidden Markov model (DNN-HMM) morgan1990continuous ; bourlard2012connectionist ; young1996review ; povey2005discriminative ; abdel2012applying ; graves2013hybrid ; dahl2011context system is typically trained, where the DNN is optimized by minimizing cross-entropy loss on the forced alignments generated for each frame of audio embeddings from a hidden Markov model-Gaussian mixture model (HMM-GMM). One notable disadvantage of the hybrid approach is that the model cannot be optimized in an E2E manner, which may result in suboptimal performance hannun2014deep . More recently, E2E models for ASR have become very popular due to their superior performance. There are three popular approaches for training an E2E model: (i) attention-based encoder-decoder (AED) models chan2015listen ; whisper ; watanabe2017hybrid ; prabhavalkar2023end , (ii) using Connectionist Temporal Classification (CTC) loss graves2006connectionist ; graves2014towards , and (iii) neural Transducer-based models graves2012sequence ; kuang2022pruned ; graves2013speech . AED models use an encoder to convert the input audio sequence into a hidden representation. The decoder, typically auto-regressive, generates the output text sequence by attending to specific parts of the input through an attention mechanism, often referred to as soft alignment mtst between the audio and text sequences. This design, however, can make it challenging to obtain word-level timestamps and to do teacher-student training with soft labels. Training AED models also requires a comparatively large amount of data, which can be prohibitive in low-resource setups. In contrast to AED models, CTC and transducer-based models maximize the marginal probability of the correct sequence of tokens (transcript) over all possible valid alignments (paths), often referred to as hard alignment mtst . However, recent research has shown that only a few paths, which are dominated by blank labels, contribute meaningfully to the marginalization, leading to the well-known peaky behavior that can result in suboptimal ASR performance ctc-peaky . Unfortunately, it is not possible to directly identify these prominent paths, or those that do not disproportionately favor blank labels, in advance within E2E models. This observation serves as the main motivation of our work.
@@
  • We propose a novel, parameterized, and differentiable seq2seq alignment model with linear complexity both in time and space.
@@
  • Finally, we conduct proof-of-concept experiments on the TIMIT garofolo1993timit , AMI ami , and Librispeech panayotov2015librispeech datasets, demonstrating that our method mitigates the peaky beahavior, improves alignment performance, and achieves promising results in E2E ASR.
@@ 2 Related Work
CTC loss. The CTC criterion graves2006connectionist is a versatile method for learning alignments between sequences. This versatility has led to its application across various seq2seq tasks st1 ; reorder ; mtst ; mt1 ; orc ; gesture . However, despite its widespread use, CTC has numerous limitations that impact its effectiveness in real-world applications. To address issues such as peaky behavior ctc-peaky , label delay tian2022bayes , and alignment drift sak2015learning , researchers have proposed various extensions. These extensions aim to refine the alignment process, ensuring better performance across diverse tasks. Delay-penalized CTC yao2023delay and blank symbol regularization yang2023blank ; zhao2022investigating ; bluche2015framewise attempt to mitigate label delay issues. Other works have tried to control alignment through teacher model spikes align1 ; align2 or external supervision externalali ; 7404851 ; 9003863 , though this increases complexity. More recently, Bayes Risk CTC tian2022bayes offer customizable, E2E approaches to improve alignment without relying on external supervision. The latest advancement, Consistency-Regularized CTC (CR-CTC) yao2024cr , mitigates extreme peaky behavior by enforcing consistency between CTC distributions obtained from different augmented views of the same audio.
Transducer loss. The transducer loss was introduced to address the conditional independence assumption of CTC by incorporating a predictor network graves2012sequence . However, similarly to CTC, transducer models suffer from label delay and peaky behavior yu2021fastemit . To mitigate these issues, several methods have been proposed, such as e.g., Pruned RNN-T kuang2022pruned , which prunes alignment paths before loss computation, FastEmit yu2021fastemit , which encourages faster symbol emission, delay-penalized transducers kang2023delay , which add a constant delay to all non-blank log-probabilities, and minimum latency training shinohara2022minimum , which augments the transducer loss with the expected latency. Further extensions include CIFTransducer for efficient alignment zhang2023say , self-alignment techniques kim2021reducing , and lightweight transducer models using CTC forced alignments wan24_interspeech .
@@ 3 Problem Formulation
We define $\mathcal{U}_{\leq N}^{d}=\bigcup_{n\leq N}\mathcal{U}_{n}^{d}$ to be the set of all $d$-dimensional vector sequences of length at most $N$. Let us consider a distribution $\mathcal{D}_{\mathcal{U}_{\leq N}^{d}\times\mathcal{U}_{\leq N}^{d}}$ and pairs of sequences $(\{{\bm{x}}_{i}\}_{i=1}^{n},\{{\bm{y}}_{i}\}_{i=1}^{m})$ of length $n$ and $m$ drawn from $\mathcal{D}_{\mathcal{U}_{\leq N}^{d}\times\mathcal{U}_{\leq N}^{d}}$. For notational simplicity, the sequences of the pairs $(\{{\bm{x}}_{i}\}_{i=1}^{n},\{{\bm{y}}_{i}\}_{i=1}^{m})$ will be respectively denoted by $\{{\bm{x}}\}_{n}$ and $\{{\bm{y}}\}_{m}$ in the following. The goal in seq2seq tasks is to train a classifier that can accurately predict the target sequence $\{{\bm{y}}\}_{m}$ from the input sequence $\{{\bm{x}}\}_{n}$, enabling it to generalize to unseen examples. Typically, $n\neq m$, creating challenges for accurate prediction as there is no natural alignment between the two sequences. In this paper, we introduce a framework to address this class of problems, applying it specifically to the ASR domain. In this context, the first sequence $\{{\bm{x}}\}_{n}$ represents an audio signal, where each vector ${\bm{x}}_{i}\in\mathbb{R}^{d}$ corresponds to a time frame in the acoustic embedding space. The second sequence $\{{\bm{y}}\}_{m}$ is the textual transcription of the audio, where each element ${\bm{y}}_{i}$ belongs to a predefined vocabulary $L=\{l_{1},\dots,l_{|L|}\}$, such that $\{{\bm{y}}\}_{m}\in L^{m}$, where $L^{m}$ denotes the set of all $m$-length sequences formed from the vocabulary $L$.
Figure 1: Alignment between embeddings of frames and target sequence. Red bullets represent the elements of the target sequence $\{{\bm{y}}\}_{m}$, while the blue bullets indicate the frame embeddings $\{{\bm{x}}\}_{n}$. In OTTC, the alignment guides the prediction model $F$ in determining which frames should map to which labels. Additionally, the alignment model has the flexibility to leave some frames unaligned, as represented by the blue-and-white bullets, allowing those frames to be dropped during inference.
@@ 4 Optimal Temporal Transport Classification
The core idea is to model the alignment between two sequences as a mapping to be learned along with the frame labels (see Figure 1). As the classification of audio frames improves, inferring the correct alignment becomes easier. Conversely, accurate alignments also improve frame classification. This mutual reinforcement between alignment and classification highlights the benefit of addressing both tasks simultaneously, contrasting with traditional hybrid models that treat them as separate tasks morgan1990continuous . To achieve this, we propose the SOTD, a framework for constructing pseudo-metrics over the sequence space $\mathcal{U}_{\leq N}^{d}$, based on a differentiable, parameterized model that learns to align sequences. Using this framework, we derive the OTTC loss, which allows the model to learn both the alignment and the classification in a unified manner.
@@ 4.2 Differentiable Temporal Alignment with Optimal Transport
In the following, we introduce 1D OT and define our alignment model. Consider the 1D discrete distributions $\mu[\bm{\alpha},n]$ and $\nu[\bm{\beta},m]$ expressed as superpositions of $\delta$ measures, i.e., a distribution that is zero everywhere except at a single point, where it integrates to 1
$\mu[\bm{\alpha},n]=\sum_{i=1}^{n}\alpha_{i}\delta_{i}\text{\>\>\>\>\>and\>\>\>\>\>}\nu[\bm{\beta},m]=\sum_{i=1}^{m}\beta_{i}\delta_{i}.$ (1)
The bins of $\mu[\bm{\alpha},n]$ and $\nu[\bm{\beta},m]$ are $\llbracket 1,n\rrbracket$ and $\llbracket 1,m\rrbracket$, respectively, whereas the weights $\alpha_{i}$ and $\beta_{i}$ are components of the vectors $\bm{\alpha}\in\Delta^{n}$ and $\bm{\beta}\in\Delta^{m}$, with $\Delta^{n}$ the simplex set defined as $\Delta^{n}=\{\mathbf{v}\in\mathbb{R}^{n}|{0\leq v_{i}\leq 1,\sum_{i=1}^{n}v_{i}=1}\}\subset\mathbb{R}^{n}$. OT theory provides an elegant and versatile framework for computing distances between distributions such as $\mu[\bm{\alpha},n]$ and $\nu[\bm{\beta},m]$, depending on the choice of the cost function peyre2019computational (chapter 2.4). One such distance is the 2-Wasserstein distance $\mathcal{W}_{2}$, which measures the minimal cost of transporting the weight of one distribution to match the other. This distance is defined as
@@
where $i-j|\!|_{2}^{2}$ is the cost of moving weight from bin $i$ to bin $j$ and $\gamma_{i,j}$ is the amount of mass moved from $i$ to $j$. The optimal coupling matrix $\bm{\gamma}^{*}$ is searched within the set of valid couplings $\Gamma^{\bm{\alpha},\bm{\beta}}$, defined as
@@
This constraint ensures that the coupling conserves mass, accurately redistributing all weights between the bins. A key property of optimal transport in 1D is its monotonicity Peyr2019NumericalOT . Specifically, if there is mass transfer between bins $i$ and $j$ (i.e., $\gamma_{i,j}^{*}>0$) and similarly between bins $k$ and $l$ (i.e., $\gamma_{k,l}^{*}>0$), then it must hold that $i\leq k\Rightarrow j\leq l$. Consequently, when $\bm{\beta}$ has no zero components – meaning that every bin from $\nu$ is reached by the transport – the set $\{(i,j)\in[\!|1,n|\!]\times[\!|1,m|\!]\ |\ \bm{\gamma}_{i,j}^{*}>0\}$ satisfies the conditions of Definition 1, thereby forming a discrete monotonic alignment. This demonstrates that the optimal coupling can effectively model such alignments (see Figure 2).
@@
Lemma 1: The function $\bm{\alpha}\mapsto\bm{\gamma}_{n}^{m,\bm{\beta}}(\bm{\alpha})$ is bijective from $\Delta^{n}$ to $\Gamma^{*,\bm{\beta}}[n]$ .
@@
4.2.1 Sequence-to-Sequence Distance